An object of mass, mi, is placed on a rough table
Answer #1

Here,

1) let the coefficient of friction is us

Normal reaction for M1 , N1 = m1 * g

Now , for the block m1 to stay stationary

m1 * g * us = m2 * g

us = m2/m1

the coeffcient of friction is m2/m1
Answer #2

The tension due to the block of mass m2 is, T m28 The force on the block of mass m, is due to the friction which is given by